A car travels for 30 minutes at 30 m/s due North. The car stops for 10 minutes, then turns and travels for 20 minutes due south at 20 m/s. What is the average velocity of the car? What is the average speed of the car?
Answer #1

distance travelled in 30 mins

d1 = speed * time = 30* 30* 60 =54000

distance travelled in 20 mins

d2 = 20* 20*60 = 24000 m

==========

a)

av velocity

v = ( 54000 - 24000) / (20 + 30) * 60 = 10 m/s

b)

av speed = ( 54000 + 24000) / ( 20 + 30)* 60 = 26 m/s

==========

do comment in case any doubt, will reply for sure . Goodluck